A 130.0 mL buffer solution is 0.105 molL?1 in NH3 and 0.135 molL?1 in NH4Br.

1.What mass of HCl will this buffer neutralize before the pH falls below 9.00?

m=____g

2.If the same volume of the buffer were 0.270 molL?1 in NH3 and 0.400 molL?1 in NH4Br, what mass of HCl could be handled before the pH fell below 9.00?

m=______g
Answer #1

Moles of NH, in 0.105 M 130 mL are as follows 130mLx0.105M 0.0136molNH, 1000mL/L Moles of NH,Brin 0.135 M, 130 mL are as foll

In a buffer solution of NH3 and NH,Br following changes occurs in the concentration of NH and NH, on addition of HCl as follo
Answer #2

NH3 is a weak base and NH4Bris its salt with a strong acid (HBr). Thus, they forma basic buffer solution. The pH of a basic bThus, using the Henderson equation: pOH 14-9 5=4.75+ log 0.105-0.21x .135+0.21x_0 5 log 0.105-0.21x 0.135+0.21x =10025 =178 02) In this case, the initial concentrations of the base and salt are different. Thus, using the Henderson equation: pOH = 14-